# Compton scattering from superstrings

### v1

@@ 4 Gravitational Compton amplitudes for all spins
$M^{\alpha\gamma,\beta\delta}_{4}({\bf 1},2,{\bf 3},4)=8i\left(\frac{(-1)^{n}}{n[\Gamma(n)]^{4}}\right)\kappa^{2}m^{2}\frac{1}{\alpha^{\prime}t(s\alpha^{\prime}-4n)(u\alpha^{\prime}-4n)}\left[{\cal A}_{0}^{\alpha\beta}{\cal A}_{0}^{\gamma\delta}\right]_{\alpha^{\prime}\to\alpha^{\prime}/4}$ (26)
where we have expanded $\sin\left(\pi x\right)$ around $x=n$ and also restored the gravity coupling constants, $M_{N}\to\left(\kappa/2\right)^{N-2}M_{N}$ [76], where $\kappa^{2}=32\pi G$. The first mass level $n=1$ corresponds in our prescription to a massive spin-2 particle since here we are compactifying two polarizations of each of the gauge amplitudes. In turn, it is easy to see that our expression agrees with the one derived in Ref. [22] in quadratic order in $a^{\mu}$, as expected from the KLT formula, see also Ref. [77]. In fact, up to quartic order in classical spin our results also match with the ones given in Ref. [22].
After imposing the three constraints listed above, the KLT relations allows us to derive our candidate opposite helicity gravity Compton amplitude which is valid for all quantum spins $S=2n>0$:
¯$\bar{\cal M}_{S}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $8i\left(\frac{(-1)^{n}}{n[\Gamma(n)]^{4}}\right)\kappa^{2}m^{2}\frac{1}{\alpha^{\prime}t(\alpha^{\prime}s-4n)(\alpha^{\prime}u-4n)}$ (27)
@@
One can also similarly construct the all plus gravitational amplitude from the corresponding electromagnetic case and the result matches the one known in the literature [69], apart from a positive multiplicative numerical factor. Moreover, the classical limit of the opposite-helicity amplitude does not satisfy the spin-shift symmetry for quantum spins greater than 2; in any case, the massive-spin 2 opposite-helicity amplitude ($S=2$) obtained from the above double copy formula preserves this symmetry.
@@
We impose the constraints on the gravitational amplitude after performing the double copy procedure on the full string amplitude. This could make the double-copy construction compatible with factorization poles that match the known 3-point amplitudes. We believe this happens because our amplitude inherits some aspects of the full original string amplitude; it is a “stringy” quantum-field-theory amplitude. To confirm this it is also necessary to exploit the R sector of string theory.
@@ 5 Conclusions and perspectives
In this work we have proposed candidate Compton amplitudes which are valid to all integer spins. An amusing feature is that the total amplitude (14) contains both Comptons for same and opposite helicities of photons (and gravitons, in the gravitational case, Eq. (26)). This observation is somewhat reminiscent of the Compton amplitude written in terms of massive spinors in the high-energy limit [68]; for instance, for the massive spin-one amplitude, the high-energy limit is given by a sum of its “helicity constituents”. As helicity amplitudes the sum of components is generally not allowed, yet this is precisely how one can sort out the different pieces of the amplitude that unifies all the different helicity amplitudes merged into a single object. Remarkably we observe a similar phenomenon in the present case.
@@
    The gravitational amplitudes obey double copy.
One possible conclusion is that subleading Regge trajectories seem to play an important role here. Indeed, as we have shown in the case of the first excited state $n=1$, the compactification down to the spin-$1$ case allowed us to recover known results in the literature and supplied the motivation to search for a valid amplitude for higher spins based on a string-theory calculation. Perhaps subleading Regge states could be the key to understand why this kind of compactification is needed here.
Our results require further explorations and many of the properties listed above are calling for a better comprehension. For instance, our soft expansion requires some refinement, perhaps following standard procedures in the literature [78, 79, 80]. On the other hand, we need to outline a better comparison with other works on the Compton amplitude such as Refs. [30, 31, 32, 33, 21, 22, 41, 58]. For instance, one could expect that contact terms may look different as in the present case they were fixed by the string excitations. These comparisons are important in order to connect our results with the existing literature and also to understand better the classical limit of our candidate Compton amplitude. Indeed, the worldsheet interpretation for Kerr provided by Ref. [15] might fit well with our description. In turn, as mentioned above, to obtain the corresponding gravitational amplitudes we not only correlate the helicites in the gauge amplitudes, but also the spins of the massive particles. So it is difficult to discuss the independence of the amplitude on the split $S\to S_{1}+S_{2}$ within our formalism. This also motivates the study of the R sector of the superstring. Such studies will be reported in a separate publication elsewhere.
@@ Appendix A. Details of the formula for the Compton scattering amplitude
${\cal A}^{\alpha\beta(1)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}$
@@
$\times$ $\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\biggr{\}}$ (35)
${\cal A}^{\alpha\beta(2)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}\left(q_{2}^{a_{S}}\eta^{\alpha c_{S}}-q_{2}^{c_{S}}\eta^{a_{S}\alpha}\right)$
@@
$+$ $\frac{(n-j)\alpha^{\prime}t}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}$ (43)
@@
-- $\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{4}^{a_{n}}\Bigr{)}\right]\biggr{\}}$ (47)
${\cal A}^{\alpha\beta(3)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}\left(q_{4}^{c_{S}}\eta^{a_{S}\beta}-q_{4}^{a_{S}}\eta^{c_{S}\beta}\right)$
@@
$+$ $\left.\frac{(n-j)\alpha^{\prime}t}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$ (55)
@@
-- $\left.\left.\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)q_{4}^{a_{n}}\Bigr{)}\right]\right\}$ (59)
@@
${\cal A}^{\alpha\beta(4)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}$
@@
-- $\left.\frac{(n-j)}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$ (67)
@@
$+$ $\left.\frac{(n-j)}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$ (73)
@@
$+$ $\left.\frac{(n-j)(n-j-1)\alpha^{\prime}t}{(2\alpha^{\prime})^{2}}\sum_{k=0}^{n-j-2}\begin{pmatrix}n-j-2\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-2-k)}q_{4}^{a(k)}\right.$ (79)
@@
$+$ $\left.\frac{(n-j)^{2}\alpha^{\prime}t}{(2\alpha^{\prime})^{2}}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$ (85)
@@
$\times$ $\left.\frac{1}{(\alpha^{\prime}t+1)}\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\right\}\eta^{a_{S}c_{S}}$ (94)
@@
and so there are no tachyons in the spectrum, as expected.
@@
${\cal S}_{\alpha^{\prime}}[i_{1},\ldots,i_{k}|j_{1},\ldots j_{k}]_{p}\equiv\left(\frac{\pi\alpha^{\prime}}{2}\right)^{-k}\prod_{t=1}^{k}\sin\left[\frac{\pi\alpha^{\prime}}{2}\left(p\cdot k_{i_{t}}+\sum_{q>t}^{k}\theta(i_{t},i_{q})k_{i_{t}}\cdot k_{i_{q}}\right)\right]$ (101)

### v2

@@ 4.1 Double copy and first mass level
$M^{\alpha\gamma,\beta\delta}_{4}({\bf 1},2,{\bf 3},4)=64i\left(\frac{(-1)^{n}}{n[\Gamma(n)]^{4}}\right)\kappa^{2}m^{2}\frac{1}{\alpha^{\prime}t(s\alpha^{\prime}-4n)(u\alpha^{\prime}-4n)}\left[{\cal A}_{0}^{\alpha\beta}{\cal A}_{0}^{\gamma\delta}\right]_{\alpha^{\prime}\to\alpha^{\prime}/4}$ (46)
where we have expanded $\sin\left(\pi x\right)$ around $x=n$ and also restored the gravity coupling constants, $M_{N}\to\kappa^{N-2}M_{N}$ [3] ($N$ is the number of external particles), where $\kappa^{2}=32\pi G$. The first mass level $n=1$ corresponds in our prescription to a massive spin-2 particle since here we are compactifying two polarizations of each of the gauge amplitudes. In turn, it is easy to see that our expression agrees with the one derived in Ref. [100] in quadratic order in $a^{\mu}$, as expected from the KLT formula, see also Ref. [151]. In fact, up to fifth order in classical spin our result for the opposite helicity amplitude also matches with the one given in Ref. [100], as we will show below. For opposite helicities, we can write that
$M_{S}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $-i\frac{\kappa^{2}}{4\,m^{2S}}\frac{\langle 4|{\bf 1}|2\bigr{]}^{4}}{(s-m^{2})(u-m^{2})t}\langle{\bf 3}|^{2S}\exp\left(i\frac{q_{4}^{\mu}\bar{\epsilon}_{4}^{\nu}{\cal J}_{1\mu\nu}}{\bar{p}\cdot\bar{\epsilon}_{4}}\right)|{\bf 1}\rangle^{2S}$ (47)
$=$ $-i\frac{\kappa^{2}}{4}\frac{1}{(s-m^{2})(u-m^{2})t}\langle 4|{\bf 1}|2\bigl{]}^{4-2S}\Bigl{(}\langle{\bf 3}4\rangle\bigl{[}{\bf 1}2\bigr{]}+\langle{\bf 1}4\rangle\bigl{[}{\bf 3}2\bigr{]}\Bigr{)}^{2S}\,\,\,\,\,(S\leq 2)$
which is just the standard result [113, 142]. As above, one can also perform the same analysis for the same-helicity amplitude, finding the usual value quoted in the literature (see the general expression below).
4.2 Double copy and all mass levels
Let us now present the gravitational Compton amplitude for even positive integer $S$. After imposing the three constraints listed above, the KLT relations allows us to derive our candidate opposite helicity gravity Compton amplitude which is valid for all quantum spins $S=2n>0$:
¯$\bar{\cal M}_{S}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $64i\left(\frac{(-1)^{n}}{n[\Gamma(n)]^{4}}\right)\kappa^{2}m^{2}\frac{1}{\alpha^{\prime}t(\alpha^{\prime}s-4n)(\alpha^{\prime}u-4n)}$ (48)
@@ 4.2 Double copy and all mass levels
One can also similarly construct the all plus gravitational amplitude from the corresponding electromagnetic case and the result matches the one known in the literature [142]
¯$\bar{\cal M}_{S}({\bf 1},2^{++},{\bf 3},4^{++})=-im^{4-2S}\frac{\kappa^{2}}{4}\frac{\bigl{[}24\bigr{]}^{4}}{t}\frac{\langle{\bf 1}{\bf 3}\rangle^{2S}}{[s-m^{2}][u-m^{2}]},$ (49)
apart from a positive multiplicative numerical factor (not shown in the expression above). This formula is valid to all $S$.
Let us see how the 3-point Kerr amplitude can also be identified as a contribution in the full closed string amplitude, after imposing the above constraints. We simply repeat the calculation for the closed superstring exploiting the simple KLT relation at three points
$M_{3}({\bf 1}_{S},{\bf 2}_{S},k)=i\kappa\left[\frac{1}{e}A_{3}({\bf 1}_{S/2},{\bf 2}_{S/2},k)\right]\left[\frac{1}{e}A_{3}({\bf 1}_{S/2},{\bf 2}_{S/2},k)\right],$ (50)
where as above we perform the rescaling $\alpha^{\prime}\to\alpha^{\prime}/4$ which is counterbalanced by the fact that the leading Regge closed string states produce $\alpha^{\prime}=2(S-2)/m^{2}$. We find that
$M_{3}({\bf 1},{\bf 2},k^{++})$ $=$ $(-1)^{S}\frac{i\kappa}{2}(mx_{1}^{+})^{2}\left(\frac{\langle{\bf 1}{\bf 2}\rangle}{m}\right)^{2S}\Bigl{(}1+{\cal O}(1/m)\Bigr{)}$
$M_{3}({\bf 1},{\bf 2},k^{--})$ $=$ $(-1)^{S}\frac{i\kappa}{2}(mx_{1}^{-})^{2}\left(\frac{\bigl{[}{\bf 1}{\bf 2}\bigr{]}}{m}\right)^{2S}\Bigl{(}1+{\cal O}(1/m)\Bigr{)}$ (51)
which of course is the usual result for the minimally coupled gravitational amplitude.
@@
Similar to the electromagnetic case, our claim here is that it is possible that closed string-theory amplitudes secretly enclose compact-body amplitudes. Furthermore, observe that we impose the constraints on the gravitational amplitude after performing the double copy procedure on the full string amplitude. This could make the double-copy construction compatible with factorization poles that match the known 3-point amplitudes. We believe this happens because our amplitude inherits some aspects of the full original string amplitude; it is a “stringy” quantum-field-theory amplitude.
4.3 Classical limit of the opposite-helicity gravitational amplitude
Here we will present the classical limit of Eq. (48). Again, as discussed above there are many works in the literature concerning the calculation of this amplitude, but for practical purposes we choose to use Ref. [107] as our main source of comparison. Many details that will not be exhibited here can be found in such a work. Again we start by rewriting the amplitude in terms of the variables $W_{\pm},U,V,\rho$ and $\chi$. We find that
¯$\bar{\cal M}_{S}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $i\left(\frac{(-1)^{n}}{[\Gamma(n)]^{4}}\right)\kappa^{2}\frac{1}{t(s-m^{2})(u-m^{2})}$ (52)
$\times$ $\left({\cal N}\frac{n^{2}(2n)^{n-1}(-1)^{n-1}}{2\,\Gamma(n)^{2}}\gamma(s\alpha^{\prime}/4,n-1)\gamma(u\alpha^{\prime}/4,n-1)\right)^{2}(\rho\cdot\chi)^{4}$
$\times$ $\left[\left(\frac{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.$
$\times$ $\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m[{\bf 1}{\bf 3}])^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}\right)\right.$
$+$ $\left.\left(\frac{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.$
$\times$ $\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m\langle{\bf 1}{\bf 3}\rangle)^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}\right)\right]^{2}$
where we recall that $S=2n$. The classical limit is obtained with
$M_{\textrm{cl}}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $\lim_{\hbar\to 0}\frac{i\,\kappa^{2}}{4}\frac{1}{t(s-m^{2})(u-m^{2})}e^{-|{\cal Z}|^{2}}(\rho\cdot\chi)^{4}$ (53)
$\times$ $\sum_{n=1}^{\infty}\frac{1}{(2(n-1))!}\left(\frac{\sqrt{\pi}}{4^{n-1}\Gamma(n)\Gamma\left(n-\frac{1}{2}\right)}\right)^{2}(-1)^{3n}$
$\times$ $\left[\left(\frac{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.$
$\times$ $\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m[{\bf 1}{\bf 3}])^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}\right)\right.$
$+$ $\left.\left(\frac{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.$
$\times$ $\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m\langle{\bf 1}{\bf 3}\rangle)^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}\right)\right]^{2}.$
To understand the $\hbar\to 0$ limit of the term $e^{-|{\cal Z}|^{2}}(\rho\cdot\chi)^{4}$, we again resort to the standard expression for $M_{S}({\bf 1},2^{++},{\bf 3},4^{--})$ with $S\leq 2$ given above. After performing similar steps as the ones carried out in the previous section, we find that
$M_{S=2,\textrm{cl}}({\bf 1},2^{++},{\bf 3},4^{--})=-i\frac{\kappa^{2}}{4}\frac{(k_{1}\cdot\chi)^{4}}{[k_{1}\cdot(q_{2}-q_{4})]^{2}(q_{2}+q_{4})^{2}}e^{x}\left(1-w+\frac{w^{2}}{2}-\frac{w^{3}}{3!}+\frac{w^{4}}{4!}\right)+{\cal O}(a^{5}).$ (54)
where again the expansion follows from the fact that we can have at most $w^{4}$ to prevent spurious poles from showing up. We can use this result in our amplitude and, as before, we prevent possible double-counting by suitably expanding $e^{x-w}$. In this way we obtain our final formula for the classical gravitational amplitude for opposite helicities:
$M_{\textrm{cl}}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $-\frac{i\,\kappa^{2}}{4}\frac{(p_{1}\cdot\chi)^{4}}{[p_{1}\cdot(q_{2}-q_{4})]^{2}(q_{2}+q_{4})^{2}}[1+(e^{x}-1)]\left[1+\left(-w+\frac{w^{2}}{2}-\frac{w^{3}}{3!}+\frac{w^{4}}{4!}\right)\right]$ (55)
$+$ $\frac{i\,\kappa^{2}}{4}\frac{(p_{1}\cdot\chi)^{4}}{[p_{1}\cdot(q_{2}-q_{4})]^{2}(q_{2}+q_{4})^{2}}\sum_{n=2}^{\infty}\frac{1}{(2(n-1))!}\left(\frac{\sqrt{\pi}}{4^{n-1}\Gamma(n)\Gamma\left(n-\frac{1}{2}\right)}\right)^{2}(-1)^{n}$
$\times$ $\left\{(e^{x}-1)\left(-w+\frac{w^{2}}{2}-\frac{w^{3}}{3!}+\frac{w^{4}}{4!}\right)\right.$
$\times$ $\left.\left[\Bigl{(}({x}+5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}+{y})^{2}-{z}^{2}}{({x}+5{y})^{2}-{z}^{2}}\right)\right.\right.$
$+$ $\left.\left.\Bigl{(}({x}-5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}-{y})^{2}-{z}^{2}}{({x}-5{y})^{2}-{z}^{2}}\right)\right]^{2}\right.$
$+$ $\left.\left(-w+\frac{w^{2}}{2}-\frac{w^{3}}{3!}+\frac{w^{4}}{4!}\right)\left[\Bigl{(}({x}+5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}+{y})^{2}-{z}^{2}}{({x}+5{y})^{2}-{z}^{2}}\right)\right.\right.$
$+$ $\left.\left.\Bigl{(}({x}-5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}-{y})^{2}-{z}^{2}}{({x}-5{y})^{2}-{z}^{2}}\right)\right]^{2}\right.$
$+$ $\left.(e^{x}-1)\left[\Bigl{(}({x}+5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}+{y})^{2}-{z}^{2}}{({x}+5{y})^{2}-{z}^{2}}\right)\right.\right.$
$+$ $\left.\left.\Bigl{(}({x}-5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}-{y})^{2}-{z}^{2}}{({x}-5{y})^{2}-{z}^{2}}\right)\right]^{2}\right\}+{\cal O}(\hbar).$
Expanding the result up to ${\cal O}(a^{5})$, our results agree with those in Ref. [107]. At higher orders, our results differ. Moreover, the classical limit of the opposite-helicity amplitude does not satisfy the spin-shift symmetry for quantum spins greater than 2; in any case, the massive-spin 2 opposite-helicity amplitude ($S=2$) obtained from the above double copy formula preserves this symmetry.
@@ 5 Conclusions and perspectives
In this work we have proposed candidate Compton amplitudes which are valid to all integer spins. An amusing feature is that the total amplitude (16) contains both Comptons for same and opposite helicities of photons (and gravitons, in the gravitational case, Eq. (46)). This observation is somewhat reminiscent of the Compton amplitude written in terms of massive spinors in the high-energy limit [141]; for instance, for the massive spin-one amplitude, the high-energy limit is given by a sum of its “helicity constituents”. As helicity amplitudes the sum of components is generally not allowed, yet this is precisely how one can sort out the different pieces of the amplitude that unifies all the different helicity amplitudes merged into a single object. Remarkably, we observe a similar phenomenon in the present case.
@@
    The gravitational quantum amplitudes obey double copy.
As we have shown in the case of the first excited state $n=1$, the compactification down to the spin-$1$ case allowed us to recover known results in the literature and supplied the motivation to search for a valid amplitude for higher spins based on a string-theory calculation. We argued that this works also for the case of the 3-point amplitude. Nevertheless, it should be noted that this is certainly an ambiguous procedure (at best) that calls for a better comprehension. One possible conclusion is that subleading Regge trajectories could be the key to understanding why this kind of exotic compactification is needed here.
While comparisons to other works could certainly be enlightening, we chose to compare our classical Compton-amplitude results to those of Refs. [106, 107], as we have emphasized above. The reason for this is that these works allow a direct parallel with our results, facilitating not only the extraction of the classical limit of the amplitudes presented here, but also making clear the physics behind the outcomes. Indeed, one can even speculate whether there is a UV completion within the framework of string theory for the massive higher-spin effective theory investigated in such references. On the other hand, our soft expansion requires some refinement, perhaps following standard procedures in the literature [153, 154, 155]. Going further, the worldsheet interpretation for Kerr provided by Ref. [96] might fit well with our description. In turn, as mentioned above, to obtain the corresponding gravitational amplitudes we not only correlate the helicites in the gauge amplitudes, but also the spins of the massive particles. So it is difficult to discuss the independence of the amplitude on the split $S\to S_{1}+S_{2}$ within our formalism. This also motivates the study of the R sector of the superstring. We intend to explore such research directions in the future.
@@ Appendix A. Details of the formula for the Compton scattering amplitude
${\cal A}^{\alpha\beta(1)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}$ (58)
@@
$\times$ $\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\biggr{\}}$
${\cal A}^{\alpha\beta(2)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}\left(q_{2}^{a_{S}}\eta^{\alpha c_{S}}-q_{2}^{c_{S}}\eta^{a_{S}\alpha}\right)$ (59)
@@
$+$ $\frac{(n-j)\alpha^{\prime}t}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}$
@@
-- $\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{4}^{a_{n}}\Bigr{)}\right]\biggr{\}}$
${\cal A}^{\alpha\beta(3)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}\left(q_{4}^{c_{S}}\eta^{a_{S}\beta}-q_{4}^{a_{S}}\eta^{c_{S}\beta}\right)$ (60)
@@
$+$ $\left.\frac{(n-j)\alpha^{\prime}t}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$
@@
-- $\left.\left.\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)q_{4}^{a_{n}}\Bigr{)}\right]\right\}$
@@
${\cal A}^{\alpha\beta(4)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}$ (61)
@@
-- $\left.\frac{(n-j)}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$
@@
$+$ $\left.\frac{(n-j)}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$
@@
$+$ $\left.\frac{(n-j)(n-j-1)\alpha^{\prime}t}{(2\alpha^{\prime})^{2}}\sum_{k=0}^{n-j-2}\begin{pmatrix}n-j-2\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-2-k)}q_{4}^{a(k)}\right.$
@@
$+$ $\left.\frac{(n-j)^{2}\alpha^{\prime}t}{(2\alpha^{\prime})^{2}}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$
@@
$\times$ $\left.\frac{1}{(\alpha^{\prime}t+1)}\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\right\}\eta^{a_{S}c_{S}}$
@@
and so there are no tachyons in the spectrum, as expected. Moreover, the coupling constant $e$ arises by demanding that the amplitude describes the scattering between massive charged particles and photons – this requirement produces $e^{-\lambda}=\frac{1}{e^{2}}$.
@@
${\cal S}_{\alpha^{\prime}}[i_{1},\ldots,i_{k}|j_{1},\ldots j_{k}]_{p}\equiv\left(\frac{\pi\alpha^{\prime}}{4}\right)^{-k}\prod_{t=1}^{k}\sin\left[\frac{\pi\alpha^{\prime}}{2}\left(p\cdot k_{i_{t}}+\sum_{q>t}^{k}\theta(i_{t},i_{q})k_{i_{t}}\cdot k_{i_{q}}\right)\right]$ (68)